# Turning Up the Heat: Min- Sampling for Creative and Coherent LLM Outputs

### v7

@@ 2 Related Work
Recent approaches like (-sampling) and mirostat sampling dynamically adjust the sampling pool based on token distribution entropy (Hewitt etΒ al., 2022a; Basu etΒ al., 2021). While these show promise, they often require complex parameter tuning and can be computationally expensive. We detail our experimental challenges running sampling in SectionΒ A.3.
@@ 3.3 Advantages Over Existing Methods
Min- is simple, requiring minimal calculations. Unlike methods involving auxiliary models (Li etΒ al., 2023a) or complex entropy-based adjustments (Hewitt etΒ al., 2022a), min- integrates easily into existing LLM inference pipelines without substantial overhead, offering practical advantages for both research and real-world applications compared to entropy-based methods such as and sampling (Hewitt etΒ al., 2022b), as we discuss in SectionΒ A.3.
@@ 3.4 Implementation Details
Choosing the Base Threshold (): Setting between 0.05 to 0.1 effectively balances creativity and coherence across tasks and models. Higher values of (e.g., 0.5-0.7 or even close to 1) can help maintain coherence and improve benchmark scores at very high temperatures, typically at the cost of diversity (see TableΒ 13).
@@
Combining with Other Techniques: While min- sampling can work with other sampling methods or repetition penalties, we recommend using it as the primary truncation method to fully leverage its dynamic capabilities. Our experiments with combined samplers (see TableΒ 12) show that multiple truncation methods can lead to double normalization issues and suboptimal performance, as each methodβs hyperparameters have so far been optimised for standalone use. We discuss potential research into combined sampling in SectionΒ D.2.
@@ 5.1 Experimental Setup
We compared min- sampling against baseline methods, including top- sampling (Holtzman etΒ al., 2020), temperature sampling, sampling (Hewitt etΒ al., 2022b), sampling (Hewitt etΒ al., 2022b) and mirostat sampling (Basu etΒ al., 2021). Main results span temperatures 0.7 and 3.0, with further tests between 0 and 5 linked in our project repository 222https://github.com/menhguin/minp_paper/. Min- and top- perform comparably at lower temperatures 0 to 0.5 (see TableΒ 10 in Appendix), with differences within error margins. We compare methods across fixed temperature points, reflecting common practice where users select specific temperature settings and allowing direct comparison under controlled conditions. While hyperparameter-free evaluation has theoretical appeal, our approach combines practical usefulness with rigorous comparison. Nonetheless, we also separately analyze tradeoffs between accuracy and diversity across hyperparameters and note Pareto frontiers (FigureΒ 2), offering a hyperparameter-agnostic view of performance efficiency.
For min-, we used base probability thresholds of and , while top- sampling employed . These hyperparameter settings were chosen based on empirical guidelines and prior research for a fair comparison (See SectionΒ A.2 for extensive discussion).
@@ Results
Extensive experiments on Llama 3 (see Appendix TableΒ 8) show min-βs benefits generalize well across various model families and scales, from 1B-70B parameters.
@@
TableΒ 2 shows min- outperforming other methods across most temperature settings. At lower temperatures , see TableΒ 10 in Appendix), min- and top- perform similarly, as decreasing temperature reduces the impact of all truncation methods by shrinking the token candidate pool. Min-βs advantage becomes more pronounced at higher temperatures, preserving problem-solving abilities even with increased diversity. Additionally, and sampling showed exponential runtime increases with temperature and failed completely at , while min- remained computationally efficient (SectionΒ A.3).
Notably, across all experiments, best scores were achieved with min- across a range of temperatures, rather than greedy decoding, challenging assumptions that greedy decoding is strictly optimal for task benchmarks (Castel etΒ al., 2022; Wiher etΒ al., 2022). See comparisons in TableΒ 10.
@@
FigureΒ 2 shows min- sampling achieves a better accuracy-creativity trade-off compared to top- sampling, consistently lying closer to the Pareto frontier and indicating more efficient performance. The greater spread of min- configurations shows its sensitivity to hyperparameter settings, allowing fine-grained control over output diversity and coherence, whereas top- configurations cluster strongly, showing less sensitivity to hyperparameter values. We further discuss this nuance in SectionΒ A.4.
@@
Independent EQ-Bench Creative Writing Paech (2024) results further support these findings, with min- (, ) scoring 62 versus the baselineβs 51.5 at (Gusev, 2024). A supplementary structured LLM-as-judge evaluation focusing on specific creative writing dimensions confirmed min-βs advantages across temperature settings (see SectionΒ C.3.6).
@@ 6 Human Evaluation
To ensure high-quality responses, we implemented several measures including attention checks and incentives for detailed feedback. For full details, see sectionΒ C.1. Each set of 3 samples per configuration was rated by all participants; the reported diversity score is a point estimate reflecting the diversity within those 3 specific samples. This setup aimed to provide valuable directional insights into perceived quality and diversity differences between methods, despite considerations limiting sample size such as retaining readersβ attention over very long reading durations, ensuring fair standardised scoring and keeping participant costs manageable.
@@ Results
Qualitative feedback provided varied perspectives on the different methods (see supplementary materials for full responses). To address limitations of the initial evaluation, we conducted a refined human study with an improved inference engine and other setup refinements (detailed in SectionΒ C.2), which reinforced our findings. Both results are available on our GitHub repository333https://github.com/menhguin/minp_paper/.
@@ 7 Conclusion
Min-βs strengths lie in its simplicity, computational efficiency, and seamless integration into existing pipelines, as evidenced by its rapid adoption in many leading open-source frameworks. By addressing the longstanding diversity-quality tradeoff in a way practitioners have found immediately useful, min- represents a notable advancement in generative language modeling, enhancing applications that require both high-quality and diverse text generation. For more detailed examples, see sectionΒ D.1.
@@
We have made demonstrable efforts to ensure the reproducibility of our results. The implementation of the proposed min- sampling method is provided in SectionΒ A.1 and is also available at our project repository.444https://github.com/menhguin/minp_paper/ Detailed descriptions of experimental setups, including model configurations, hyperparameter settings, and evaluation protocols, are outlined in SectionΒ 5 and SectionΒ A.2. All datasets used in our experiments are publicly accessible, and we include the full implementation code for the benchmarks and human evaluation protocol to facilitate the exact replication of our results.
@@ 6. Quality Control Measures
Two strategically placed attention checks were embedded in the survey at the 40%Β and 70%Β marks:
@@ D.2 Limitations and Future Research Directions
The base probability threshold is a critical hyperparameter. Investigating methods for dynamically adjusting based on context or developing guidelines for optimal settings across various tasks could enhance performance. This was particularly challenging, as MAUVE hyperparameter sweeps are measured without temperature scaling on GPT2-XL (Pillutla etΒ al., 2023), and require pre-selection on exact hyperparameters. min-, however, is a novel method often used in conjunction with temperature scaling without extensive literature/experimental data. We discuss this further in SectionΒ A.2.

### v8

@@ 2 Related Work
Recent approaches like (-sampling) and mirostat sampling dynamically adjust the sampling pool based on token distribution entropy (Hewitt etΒ al., 2022a; Basu etΒ al., 2021). While these show promise, they often require complex parameter tuning and can be computationally expensive. We detail our experimental challenges running sampling in SectionΛA.3.
@@ 3.3 Advantages Over Existing Methods
Min- is simple, requiring minimal calculations. Unlike methods involving auxiliary models (Li etΒ al., 2023a) or complex entropy-based adjustments (Hewitt etΒ al., 2022a), min- integrates easily into existing LLM inference pipelines without substantial overhead, offering practical advantages for both research and real-world applications compared to entropy-based methods such as and sampling (Hewitt etΒ al., 2022b), as we discuss in SectionΛA.3.
@@ 3.4 Implementation Details
Choosing the Base Threshold (): Setting between 0.05 to 0.1 effectively balances creativity and coherence across tasks and models. Higher values of (e.g., 0.5-0.7 or even close to 1) can help maintain coherence and improve benchmark scores at very high temperatures, typically at the cost of diversity (see TableΛ13).
@@
Combining with Other Techniques: While min- sampling can work with other sampling methods or repetition penalties, we recommend using it as the primary truncation method to fully leverage its dynamic capabilities. Our experiments with combined samplers (see TableΛ12) show that multiple truncation methods can lead to double normalization issues and suboptimal performance, as each methodβs hyperparameters have so far been optimised for standalone use. We discuss potential research into combined sampling in SectionΛD.2.
@@ 5.1 Experimental Setup
We compared min- sampling against baseline methods, including top- sampling (Holtzman etΒ al., 2020), temperature sampling, sampling (Hewitt etΒ al., 2022b), sampling (Hewitt etΒ al., 2022b) and mirostat sampling (Basu etΒ al., 2021). Main results span temperatures 0.7 and 3.0, with further tests between 0 and 5 linked in our project repository 222https://github.com/menhguin/minp_paper/. Min- and top- perform comparably at lower temperatures 0 to 0.5 (see TableΛ10 in Appendix), with differences within error margins. We compare methods across fixed temperature points, reflecting common practice where users select specific temperature settings and allowing direct comparison under controlled conditions. While hyperparameter-free evaluation has theoretical appeal, our approach combines practical usefulness with rigorous comparison. Nonetheless, we also separately analyze tradeoffs between accuracy and diversity across hyperparameters and note Pareto frontiers (FigureΒ 2), offering a hyperparameter-agnostic view of performance efficiency.
For min-, we used base probability thresholds of and , while top- sampling employed . These hyperparameter settings were chosen based on empirical guidelines and prior research for a fair comparison (See SectionΛA.2 for extensive discussion).
@@ Results
Extensive experiments on Llama 3 (see Appendix TableΛ8) show min-βs benefits generalize well across various model families and scales, from 1B-70B parameters.
@@
TableΒ 2 shows min- outperforming other methods across most temperature settings. At lower temperatures , see TableΛ10 in Appendix), min- and top- perform similarly, as decreasing temperature reduces the impact of all truncation methods by shrinking the token candidate pool. Min-βs advantage becomes more pronounced at higher temperatures, preserving problem-solving abilities even with increased diversity. Additionally, and sampling showed exponential runtime increases with temperature and failed completely at , while min- remained computationally efficient (SectionΛA.3).
Notably, across all experiments, best scores were achieved with min- across a range of temperatures, rather than greedy decoding, challenging assumptions that greedy decoding is strictly optimal for task benchmarks (Castel etΒ al., 2022; Wiher etΒ al., 2022). See comparisons in TableΛ10.
@@
FigureΒ 2 shows min- sampling achieves a better accuracy-creativity trade-off compared to top- sampling, consistently lying closer to the Pareto frontier and indicating more efficient performance. The greater spread of min- configurations shows its sensitivity to hyperparameter settings, allowing fine-grained control over output diversity and coherence, whereas top- configurations cluster strongly, showing less sensitivity to hyperparameter values. We further discuss this nuance in SectionΛA.4.
@@
Independent EQ-Bench Creative Writing Paech (2024) results further support these findings, with min- (, ) scoring 62 versus the baselineβs 51.5 at (Gusev, 2024). A supplementary structured LLM-as-judge evaluation focusing on specific creative writing dimensions confirmed min-βs advantages across temperature settings (see SectionΛC.3.6).
@@ 6 Human Evaluation
To ensure high-quality responses, we implemented several measures including attention checks and incentives for detailed feedback. For full details, see sectionΛC.1. Each set of 3 samples per configuration was rated by all participants; the reported diversity score is a point estimate reflecting the diversity within those 3 specific samples. This setup aimed to provide valuable directional insights into perceived quality and diversity differences between methods, despite considerations limiting sample size such as retaining readersβ attention over very long reading durations, ensuring fair standardised scoring and keeping participant costs manageable.
@@ Results
Qualitative feedback provided varied perspectives on the different methods (see supplementary materials for full responses). To address limitations of the initial evaluation, we conducted a refined human study with an improved inference engine and other setup refinements (detailed in SectionΛC.2), which reinforced our findings. Both results are available on our GitHub repository333https://github.com/menhguin/minp_paper/.
@@ 7 Conclusion
Min-βs strengths lie in its simplicity, computational efficiency, and seamless integration into existing pipelines, as evidenced by its rapid adoption in many leading open-source frameworks. By addressing the longstanding diversity-quality tradeoff in a way practitioners have found immediately useful, min- represents a notable advancement in generative language modeling, enhancing applications that require both high-quality and diverse text generation. For more detailed examples, see sectionΛD.1.
@@
We have made demonstrable efforts to ensure the reproducibility of our results. The implementation of the proposed min- sampling method is provided in SectionΛA.1 and is also available at our project repository.444https://github.com/menhguin/minp_paper/ Detailed descriptions of experimental setups, including model configurations, hyperparameter settings, and evaluation protocols, are outlined in SectionΛ5 and SectionΛA.2. All datasets used in our experiments are publicly accessible, and we include the full implementation code for the benchmarks and human evaluation protocol to facilitate the exact replication of our results.
@@ 6. Quality Control Measures
Two strategically placed attention checks were embedded in the survey at the 40% and 70% marks:
@@ D.2 Limitations and Future Research Directions
The base probability threshold is a critical hyperparameter. Investigating methods for dynamically adjusting based on context or developing guidelines for optimal settings across various tasks could enhance performance. This was particularly challenging, as MAUVE hyperparameter sweeps are measured without temperature scaling on GPT2-XL (Pillutla etΒ al., 2023), and require pre-selection on exact hyperparameters. min-, however, is a novel method often used in conjunction with temperature scaling without extensive literature/experimental data. We discuss this further in SectionΛA.2.